Find the roots of the polynomial equation.
-x3 + 5x2 - 11x + 55 = 0
A) i√11, -1√11,-5
B) √11,-√11,5
C) i√11,-5
D) i√11,-i√11,5

Found 2 solutions by lynnlo, ewatrrr:
Answer by lynnlo(4176)
Answer by ewatrrr(24785)

 
Hi,
y = -x3 + 5x2 - 11x + 55 ||x = 5 one root -125 + 125 - 55 + 55 = 0
using Synthetic Divsion
5 -1 5 -11 55
-5 0 -55
-1 0 -11 0
results in:
-x^2 - 11 = 0
x^2 + 11 = 0
x^2 = -11
x = ± i√11
roots are: i√11,-i√11,5